OpenStudy (anonymous):

use 1.7 as th value for sqrt3 wat is the area of equilateral triangle abc

OpenStudy (raden):

what is the length of sides that triangle ?

OpenStudy (anonymous):

8

OpenStudy (raden):

ok, use this formula : area = 1/4 * s^2 sqrt(3) (only for equilateral triangle) with s is the side of triangle

OpenStudy (anonymous):

no i gotta use 1/2bh

OpenStudy (raden):

that would be same

OpenStudy (anonymous):

oh idk how to do it yur way tho

OpenStudy (raden):

just subt s=8 area = 1/4 * 8^2 sqrt(3) = 1/4 * 64 * 1.7 = ...

OpenStudy (anonymous):

ok kool i like tht way better thnx

OpenStudy (raden):

welcome ;)

OpenStudy (anonymous):

can yu help me more?

OpenStudy (raden):

sure, if i can

OpenStudy (anonymous):

wats th formula for area of a trapezoid

OpenStudy (raden):

area = 1/2 * (b1 + b2)h with b are the bases, and h its height

OpenStudy (raden):



OpenStudy (anonymous):